Downstairs (non-Maser) Instrumentation Log

as of Sunday, 17-Dec-2017 17:33:20 AST

Go to Most Recent Entry


NameEmailDateComment
lswlwray@naic.eduTuesday, March 15, 2005 14:27:51Log Page created.
ansanolla@naic.eduFriday, May 20, 2005 12:49:27The caldipole HS had the sync pulse bad due that during greasing the ballbearings at the antenna they broke the reed relay.
W. Genterwgenter@naic.eduMonday, June 27, 2005 11:16:044-us matched filters repaired. There was a defective power supply which made the chassis appear to be osscillating.
ansanolla@naic.eduWednesday, July 6, 2005 09:59:53There were two burned pilot ligths at the cable car disable panel. One was the cable car disable indicator and the other one the SETI recvr on/off sw.The bulbs are in the stock room , they are 12 VDC 100ma EBT ones.
ansanolla@naic.eduMonday, July 11, 2005 11:55:29The remote cryo temp reading is not working due to a bad fiber optic cable.
ansanolla@naic.eduWednesday, July 13, 2005 08:46:53Phil & Jorge still working with ethernet system in order to return the remote cryo temp readings.
ansanolla@naic.eduFriday, July 15, 2005 11:19:12System back to normal. JRV & CR installed a new ethernet unit at the Dome.
ansanolla@naic.eduFriday, July 22, 2005 16:10:24There is a new 2380 Mhz SB xmitter Doppler driver at the receiver room.It is wider than the previous one.Jon Luke will run his SB run on 7/23/2005 using it.
ansanolla@naic.eduThursday, July 28, 2005 13:09:20The BNC input connector for the multiplier drive was loose.Also the HS ant indicator (reed sw) was not working due that the reed sw was removed from normal position.Reset it.
ansanolla@naic.eduFriday, August 5, 2005 10:58:23The Jon Luke (NEW) SB xmitter dopler driver was replaced by the normal one (OLD).
ansanolla@naic.eduFriday, August 5, 2005 11:27:59The old xmitter doppler was not working because they didn't swap the software program from Jon Luke to Normal one.
ansanolla@naic.eduMonday, August 29, 2005 10:48:24Test
ansanolla@naic.eduWednesday, August 31, 2005 16:14:34The HP meter 34970 used to measure the receivers cryo temperatures is not working.Has to be replaced ASAP , no Dewar cryo readings downstairs.
ansanolla@naic.eduThursday, September 1, 2005 11:30:58HP meter 34970 replacement installed , it reads the Dewar cryo temps.
ansanolla@naic.eduTuesday, September 13, 2005 15:49:24There is a problem with the 2 Ghz converter.It looks like if the L.O. amp is dead.
ansanolla@naic.eduWednesday, October 12, 2005 10:24:02The Hi side cal dipole rotating antenna motor burned out.Sent motor for rewiring.
ansanolla@naic.eduFriday, October 14, 2005 11:33:35The Dome old air conditioner at the Dome unit was producing rfi at 327.5 Mhz. So in order to solve the problem LM & JRB installed a ckt. to turn off the main power of these AC unit from downstairs.The ON/OFF sw is next to the AC remote control chassis in the radar room.
W. Genterwgenter@naic.eduWednesday, December 14, 2005 08:56:30Reported problem when using SB-RADAR. Found 30MHz Baseband mixer in rack 8-10 with no output.
ansanolla@naic.eduWednesday, December 14, 2005 13:53:10The 20-40 Mhz Baseband mixer(chassis 8-10) in rack 8 had its +/- 15 vdc power supply bad ,replaced with a new one. Also an LED power indicator added as its fuse holder.
ansanolla@naic.eduWednesday, December 28, 2005 10:15:20Both 20/40 Mhz Base band mixers at the Radar Rack 7 & 8 at Radar Room( R-685 A & B )were repaired and adjusted.Ganesh adjusted its dc offset that were out.On channel B of R-685A he found that the Sin Op amp board had a burned out 50 ohm resistor.
ansanolla@naic.eduTuesday, January 17, 2006 13:12:35The power supply for the Gaussian & Butterworth filters 8-6 & 8-8 was removed from its location in order to avoid overheating of the BB mixer 8-10.It was next to it blocking its ventilation.Lower it down.
ansanolla@naic.eduWednesday, January 18, 2006 15:15:27Radar room cabinets cooling fan air filters were replaced.All were cloged with dirt.
ansanolla@naic.eduThursday, January 19, 2006 16:19:00The 30 Mhz Butterworth filter ( 1 Mhz BW ) chassis 7-8 was dead. It was due that its output connector had a broken wire.
ansanolla@naic.eduTuesday, January 31, 2006 13:59:11The knob of 10 db variable attenuators (chassis 7-5 & 8-5,1 db step) bank in rack 7 & 8 were misplaced giving a wrong reading.
ansanolla@naic.eduTuesday, January 31, 2006 14:02:34The IF/LO sw for LO 1 was bad ,replaced with the spare one.
ansanolla@naic.eduThursday, February 2, 2006 09:48:21The Cal Dipoles antennas ( Low & Hi ) had several problems.The Low side had the sync pulse bad due to a broken reed relay , replaced and is ok. Also there was not AC power for the multiplier and was that the control relay (W88CPX-17) was bad, replaced and is ok. At the high side a similar problem occured , no sync pulse due to broken reed relay. Also the multiplier power supply was bad as all the rf amps of the unit , all were burned out,all replaced.
ansanolla@naic.eduTuesday, February 7, 2006 09:20:29BG work out with the RM-852 baseband mixers at the visitor room.He measured the levels on the unit plus check its components specifications on it.
W. Genterwgenter@naic.eduMonday, February 20, 2006 10:55:34Reported connector problems with 30MHz IF Distribution Amplifiers in Rack 8-4. We found that the screw holding the input amplifier had stripped. We re-tapped the bracket and now the input connector is OK. Upon installation we found the cooling fan to be inoperable and the fuse holder was broken. We replaced both.The chassis has been installed and operational.
ansanolla@naic.eduMonday, March 6, 2006 12:55:45The 4 usec match filter on radar room rack 8 had a bad power supply on it.Jose Rosa replaced it.
ansanolla@naic.eduTuesday, May 9, 2006 14:43:14The heliax antennas for polarity check were repaired and recalibrated.
ansanolla@naic.eduTuesday, May 16, 2006 11:51:16Forgot to specifie the frecuency of the antennas on the last log entry.Antennas are for 430 mhz.
ansanolla@naic.eduTuesday, May 16, 2006 11:57:29The rf coax cables (14 of them) for patching the GALFA cabinet at the visitors room with the SERENDIP one were installed.At the SERENDIP side a hole must be made at the floor in order to pass all these cables into the cabinet.
ansanolla@naic.eduTuesday, May 23, 2006 15:04:34The VLBI BBM #5 installed.Plug in connector drive screws mounted backwards.
ansanolla@naic.eduTuesday, May 30, 2006 15:57:08The ALFA IF amp att. at the WAPP room was reset with fixed att in order to have an output of -30 dbm (+/- 3 db)per channel (0 to 6).
ansanolla@naic.eduWednesday, May 31, 2006 15:27:23SERENDIP rack removed from Visitors room.New SERENDIP system been installed.
ansanolla@naic.eduThursday, August 3, 2006 13:46:06The radar room 30 mhz IF amp 8-3,was having problems with the 10 db gain waffwer sw. Waffers were replaced at the sw.
W. Genterwgenter@naic.eduMonday, August 7, 2006 16:22:12The left hand Krohn-Hite filter in rack 9 was reported as not operational. The DC fuse was blown. The fuse was replaced and the unit is operational.
ansanolla@naic.eduTuesday, August 8, 2006 13:29:05The Kron-Hite filter at the radar room blew the fuse again.So JV swap the ac fuse holder in order to hold the dc fuse , so is going to be easy and faster to replace it.
ansanolla@naic.eduTuesday, August 8, 2006 14:32:45JR filled up a work order to rebuilt the Helium tanks rack at the werehouse area.The rack was not identify properly causing problems to the suppliers and our W.H. personnel.
Jesus RiosThursday, August 31, 2006 10:01:34Wednesday 30 August 2006, LO Distribution chassis on rack 5-7 was removed to replace a bad HP Coaxial Switch(SP4T) model 87104A at 4:40pm.The switch location is on SW3 and the new one is Agilent serial # MY42147499 same model.
ansanolla@naic.eduWednesday, September 27, 2006 12:02:17A reset sw was added to the radar blanker. Also during its installation the switching 5vdc power supply quit working , replaced it.
anollaanolla@naic.eduThursday, October 26, 2006 16:05:29The cal antenna had a few bad things like the power indicators for both sides.Also both sync reed sw's were bad.Still have to check the rf part.
ansanolla@naic.eduWednesday, November 1, 2006 09:31:11Both multipliers were bad . The low side had a Miteq amp burned out & its pin diode attenuator ( 3 diodes bad ).Also the high side had a burned first stage of its input amp . And its pin diode att. also was burned.
Billwgenter@naic.eduMonday, January 8, 2007 14:02:46The Left hand Kron-Hite filter at the radar room blew the fuse again. Replaced fuse and informed operator and Jose Cruz about location of fuse.
ansanolla@naic.eduWednesday, February 21, 2007 16:11:37The high side Cal dipole is not radiating. Ganesg tried to use it with the LBW and didn't see nothing.
ansanolla@naic.eduTuesday, February 27, 2007 15:04:26The input amp not working , cold solder joint at output of it.Also the attenuator pot at the radar room was open.All In/Out levels taken with the HP Spec Anal. 8591A.
ansanolla@naic.eduTuesday, March 6, 2007 10:10:58The ground keeping personnel is taking care of the Cal Dipole facility , painting and cleaning the structure.Also both systems (High & Low sides) were checked by electronic failures and were fixed.Working in a new manual for it , as always its file number is R-359.
ansanolla@naic.eduTuesday, March 6, 2007 10:32:10Yesterday (3/5/2007) the .5 & 4 usec match filter bank was checked due to a complaint about the 4 usec ones.Were checked and were ok even the .5 usec ones.We check this morning with the users of the system about its performance and they said that everything was working fine ( Jean Luke & M.N.)
dana whitlowdwhitlow@naic.eduMonday, February 4, 2008 09:58:15Inserted IF splitter for mock spectrometers in signal path to WAPPS. Change was made approximately mid November 2007.
crvTuesday, April 15, 2008 16:01:03Alfa IF Attenuator switch chassis channel 0B had a 6db attenuator replaced for a 3db.
crvFriday, May 9, 2008 15:14:28Square law detector power supply replaced because it had ripples in the -15v output.
Dana Whitlowdwhitlow@naic.eduTuesday, October 14, 2008 07:48:20As of Friday 10 Oct 2008 all 14 PDEV units have had their ADC input ranges set to the more sensitive range (768 mV pp). Half the units (group 0, aka 'PALFA') had this change performed earlier in the year (April 2008).
William Iguinawiguina@naic.eduMonday, January 11, 2010 15:44:40Remplazé Dc to DC converter del Halfband Filter del wapp #4 (48VC5C100A)1/08/10
OCocolon@naic.eduFriday, April 9, 2010 08:08:35As per Ast Time:04:14 Project :Info Obs Type: Comments:Observer is reporting problems with wapp 4, can't adjust power levels. Se verifico wapp 4, el halfband estaba apagado, tenia el fusible fundido.
crvcrios@naic.eduThursday, July 22, 2010 14:57:13On Monday July 19, 2010 the transfer switch chassis R1321 on rack 5, position 4 was removed and opened due to a ripple on CH B. Several connectors were loose. After re tightening the connectors the transfer switch chassis was installed. The ripple on CH B is gone.


Back to Top